Advertising firm Harry Davis Industries has asked you to estimate its weighted average cost of capital. To that end, they have provided you with the following information:

  •  Tax rate is 40%

  •  Harry Davis can issue bonds with a 12% semiannual coupon, a $1000 par value, and a 15 year

    maturity, at a price of $1153.72 net of floatation costs.

  •  The firm has no preferred stock.

  •  The firm’s common stock is currently selling at a price of $50 per share. The most recent dividend

    was $3.12, and dividends are expected to grow at a constant annual rate of 5.8%. Harry Davis’s beta is 1.2, the yield on treasury securities is 5.6%, and the market risk premium is 6%. For the bond-yield- plus-risk-premium method, Harry Davis uses a company risk premium of 3.2%.

  •  The firm’s target capital structure is 30% debt, 70% common equity.

  1. Compute the firm’s after-tax cost of debt.

  2. Compute the firm’s cost of equity.

  3. Compute the firm’s weighted average cost of capital.

  4. Harry Davis is interested in establishing a new division focused on developing web-based content. This

    division is expected to have a beta of 1.7. It will have the same capital structure as the reset of the firm, and debt will have the same cost as for the rest of the firm. Estimate the cost of capital for this new division.
Answer #1

A)Using financial calculator to calculate the ytm of bond

Inputs: N= 15 × 2 = 30

Fv=1,000

Pv= -1,153.72

Pmt= 12% / 2 = 6% × 1,000 = 60

I/y = compute

We get, ytm of the bond as 5% × 2 = 10%

After tax cost of debt= ytm (1 - tax rate)

= 10% (1 - 0.40)

= 10% (0.60)

= 6%

After tax cost of debt is 6%

B) Cost of equity

Using CAPM model

Expected return= Risk free rate + 2 × (market risk premium)

= 5.6% + 1.2 (6%)

= 5.6% + 7.2%

= 12.8%

Using bond yield plus risk premium

= ytm of the bond + risk premium

= 10% + 3.2%

= 13.2%

Using DDM model

Cost of equity = Dividend paid ( 1+ growth rate) / price + growth rate

= 3.12 ( 1+ 0.058) / 50 + 0.058

= 3.3010 / 50 + 0.058

= 0.0660 + 0.058

= 0.1240 or 12.40%

C) Wacc= weight of debt × cost of debt + weight of equity × cost of equity

= 0.30 × 6% + 0.70 × 12.8%

= 1.8% + 8.96%

= 10.76%

D) New beta = 1.7

Using CAPM

Expected return= Risk free rate + beta ( market risk premium)

= 5.6% + 1.7 (6%)

= 5.6% + 9.52%

= 15.12%

Cost of capital for new firm = weight of debt × cost of debt + weight of equity × cost of equity

= 0.30 ×6% + 0.70 × 15.12%

= 1.8% + 10.584%

= 12.384%